ABC Sports, a store that sells various types of sports clothing and other sports items, is planning to introduce a new design of Arizona Diamondbacks’ baseball caps. A consultant has estimated the demand curve to be

Q = 2,000 - 100P

where Q is cap sales and P is price.

a. How many caps could ABC sell at $6 each?

b. How much would the price have to be to sell 1,800 caps?

c. Suppose ABC were to use the caps as a promotion. How many caps could ABC give away free?

d. At what price would no caps be sold?

e. Calculate the point price elasticity of demand at a price of $6.


1
Expert's answer
2020-10-13T07:14:55-0400

a) "\bold {Answer}"

"Q = 1,400 \space Caps"


"\bold {Solution}"

When P = $6, "Q = 2,000 - 100(6)"

"= 2,000 - 600"

"= 1,400 \space caps"



b) "\bold {Answer}"

"P = \$2 \space per \space Cap"


"\bold {Solution}"

When "Q = 1,800 \space Caps"

"=> 1,800 = 2,000 - 100P"

"=> 100P = 2,000 - 1,800"

"=> 100P = 200"

"P = \dfrac {200}{100}"

"P = \$2"


c) "\bold {Answer}"

"Q = 2,000 \space Caps"


"\bold {Solution}"

We need Q when P = $0

Thus, "Q = 2,000 - 100(\$0)"

"= 2,000 - 0"

"= 2,000 \space Caps"



d) "\bold {Answer}"

"P = \$20 \space per \space Cap"


"\bold {Solution}"

We need P when Q = 0 caps,

"Thus, \space 0 = 2,000 - 100P"

"=> 100P = 2,000"

"=> P = \dfrac {2,000}{100}"


"\therefore \space P = \$20" per cap



e) "\bold {Answer}"


"\eta = -0.429"


"\bold {Solution}"

"\eta = \dfrac {dQ}{dP} × \dfrac {P_{0}}{Q_{0}}"


When "P = \$6, \space Q = 1,400 \space Caps"


"\dfrac {dQ}{dP} = \dfrac {d}{dP} (2,000-100P)"


"= -100"


"=> \eta = \dfrac {-100}{1} × \dfrac {6}{1,400}"


"= \dfrac {-1}{1}×\dfrac {6}{14}"


"= -0.4285714285"


"= -0.429"